OpenStudy (ria23):

Solve for x 2=log(5x+12)

OpenStudy (asnaseer):

make use of the fact that if:\[\log_ab=c\]then this implies that:\[b=a^c\]

OpenStudy (ria23):

I did one similar to this one, it had a log base of 3... That became my exponent~ I don't have a base... That automatically means it's 10 right?

hartnn (hartnn):

correct :)

OpenStudy (ria23):

So I would set it up \[5x+12=10^{2}\]?

OpenStudy (asnaseer):

perfect! :)

OpenStudy (ria23):

Would that take me to \[5x-88=0 \] then factor 88 and find x from there?

hartnn (hartnn):

factoring 88 might not be useful add 88 on both sides first

OpenStudy (asnaseer):

then divide both sides by 5 and you are done :)

OpenStudy (ria23):

So I subtract 100 from both sides then add 88 on both?

hartnn (hartnn):

yes. 5x+12 = 100 5x +12-100 = 100 -100 5x-88 = 0 5x -88 + 88 = 88 5x = 88

OpenStudy (ria23):

I got 17.6 I'm not supposed to use mixed numbers, or decimal approximations of irrational numbers... Would this be the answer, or would I have to change it?

hartnn (hartnn):

isn't a fraction allowed? 5x =88 x= 88/5

OpenStudy (ria23):

>~< Oh yea. xD Thank yhu both so much for yhur help~

OpenStudy (ria23):

Would yhu mind helping me with one more?

hartnn (hartnn):

welcome ^_^ and sure, i'd be glad to :)

OpenStudy (ria23):

Alright n_n I'll open up a new thread and tag yhu. Thank yhu again! I really appreciate it~

OpenStudy (solomonzelman):

I was wondering, why not yes. 5x+12 = 100 5x = 100 -12 5x = 88 x = 88/5 lol